you land a job as a laboratory assistant over the summer break. On your second day of your job, the laboratory technician asks you to prepare a 750 mL solution of potassium permanganate (KMnO4) that is 0.489 mol/L. What mass must you collect in order to prepare the solution?
1
Expert's answer
2020-06-02T14:08:18-0400

c=n/V

n=V*c=0.750*0.489=0.367mol


M(KMnO4)=39+55+16*4=158g/mol

m=n*M=0.367*158=57.97g


Answer: m=57.97g